OpenStudy (anonymous):

in one U.S city the taxi cost is $5 plus $0.70 per mile. If you are traveling from the airport, there is an additional charge of $3.00 for tolls. How far can you travel from the airport by taxi for $32.50?

OpenStudy (anonymous):

Equaltion: 5+3+0.7(x)=32.5 0.7x= 32.5-8 =24.5 X= 24.5/0.7=35miles

OpenStudy (anonymous):

Awesome thanks :)